OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 5, -3, and -1 + 3i

OpenStudy (anonymous):

f(x) = x4 + 12.5x2 - 50x - 150 f(x) = x4 - 9x2 - 50x - 150 f(x) = x4 - 4x3 + 15x2 + 25x + 150 f(x) = x4 - 4x3 - 15x2 - 25x - 150

OpenStudy (anonymous):

I got C but not sure if im correct!

OpenStudy (campbell_st):

well look at the complex root x = -1 + 3i then x + 1 = 3i square both sides \[(x + 1)^2 = -9\] so \[x^2 + 2x + 10 = 0\] which means a quadratic factor is \[x^2 + 2x + 10\] you know the real roots x = 5 so (x - 5) is a factor x = -3 so (x + 3) is a factor then the polynomial is \[P(x) = (x^2 + 2x + 10)(x -5)(x + 3)\] so distribute to get the answer and I don't think it's C

OpenStudy (anonymous):

f(x) = x4 - 9x2 - 50x - 150?

OpenStudy (campbell_st):

that's what I got...

OpenStudy (anonymous):

After I distributed thats what i got!

OpenStudy (anonymous):

oh okay :) thanks @campbell_st

OpenStudy (campbell_st):

since we agreee it looks like A might be the answer

OpenStudy (anonymous):

yes :) @campbell_st

OpenStudy (campbell_st):

great, well done